A candy store finds that it can make a profit of P dollars each month by selling x boxes of candy. Using the formula: P(x) = -.0013x2 + 5.5x – 800, how many boxes of candy must the store sell in order to maximize their profits? What is the maximum profit?
Answer by stanbon(75887)

A candy store finds that it can make a profit of P dollars each month by selling x boxes of candy. Using the formula: P(x) = -.0013x2 + 5.5x – 800, how many boxes of candy must the store sell in order to maximize their profits? What is the maximum profit?
---
Max occurs when x = -b/(2a) = -5.5/(2*-0.0013) = 2115.38
-----
Max profit = P(2115.38) = $5017.30
--------
Cheers,
Stan H.
-----------